# Interpreting Context Look-ups in Transformers: Investigating Attention-MLP Interactions

### v1

@@ Appendix C GPT-4 Prompts
We are studying attention heads in a transformer architecture neural network. Each attention head looks for some particular thing in a short document. This attention head in particular helps to predict that the last token is “ as”, but it is only active in some documents and not others. Look at the documents and explain what makes the attention head active, taking into consideration the inactive examples. Examples where the attention head is active: could begin as early as will be available in schools as early as The lead would reach as high as Examples where the attention head is inactive: a video game console such as whatever he wants so long as Explanation: This attention head is active when the document
Figure 8: GPT-4 prompt for generating an explanation for attention head activity based on max-activating prompts it is active and inactive for. This prompt was written in a similar style to the neuron-explanation prompt from Foote et al. (2023). The prompt’s examples were truncated for readability.
Is the given example an active example? (Yes/No) This head is active when the document contains a phrase that describes a specific condition, purpose or limitation before the word “only.” In active examples, we see phrases like "for informational purposes only," "for one reason and one reason only," "for general information purposes only," or "for entertainment purposes only." These phrases define a specific context or reason which leads to the appearance of the word "only" at the end. On the other hand, in inactive examples, the word "only" is used in different contexts, without a clear specific condition or purpose preceding it.. Example: "All graphs, charts and other visual representations are shown for illustrative purposes only…" Answer:
@@
Examples where the attention head is active: out huge amounts of money to go there … of people willing to shell out huge amounts of money to go there How many minutes are there between 4:18 PM … How many minutes are there accustomed - is there … time to become accustomed - is there Explanation: The head is active when the document
contains a more complex context or multiple ideas and events...
@@ Appendix E Sample of Explainable Attention Head
Explainable Attention Heads When an attention head has a good explanation, it can be used to predict whether the attention head activates its associated neuron for a given prompt. For example, for an attention head associated with a ’ way’ neuron, if we have the following explanation: Explanation. The attention head in question seems to activate when the phrase "all the way" is used to indicate a complete path, journey, or exhaustive explanation. It is likely looking for contexts where "way" is at the end of a phrase that describes a comprehensive or continuous extent of something. In contrast, the attention head is inactive when "way" is used within idiomatic expressions, comparative phrases (e.g., "the other way around"), or as part of a more metaphorical or abstract concept (e.g., "think of it that way," "put it this way," "stay that way"). The autolabeller successfully classifies the following examples: Examples where the head is active: FROM 0 TO 5, AND THE Y AXIS MUST INCLUDE THE VALUES FROM 236 ALL THE way stretches as far west as South Dakota, through the northernmost reaches of the U.S. soybean-producing region, all the way bounding from the Oklahoma panhandle in the 1870s all the way Examples where the head is inactive: about why things are the way know G-d’s plans/why He arranges things the way "Wait a minute, it does not work that way why things are the way crosses under the New York State Thruway (Interstate 87). Past the Thruway, the commercial buildings give way

### v2

@@ Appendix C GPT-4 Prompts
GPT-4 Prompt for Explanation We are studying attention heads in a transformer architecture neural network. Each attention head looks for some particular thing in a short document. This attention head in particular helps to predict that the last token is “ as”, but it is only active in some documents and not others. Look at the documents and explain what makes the attention head active, taking into consideration the inactive examples. Examples where the attention head is active: could begin as early as will be available in schools as early as The lead would reach as high as Examples where the attention head is inactive: a video game console such as whatever he wants so long as Explanation: This attention head is active when the document
Figure 8: GPT-4 prompt for generating an explanation for attention head activity based on max-activating prompts it is active and inactive for. This prompt was written in a similar style to the neuron-explanation prompt from Foote et al. (2023). The number of examples in this prompt were shortened for readability (There are typically 15-20 examples total).
GPT-4 Prompt for Classification Is the given example an active example? (Yes/No) This head is active when the document contains a phrase that describes a specific condition, purpose or limitation before the word “only.” In active examples, we see phrases like "for informational purposes only," "for one reason and one reason only," "for general information purposes only," or "for entertainment purposes only." These phrases define a specific context or reason which leads to the appearance of the word "only" at the end. On the other hand, in inactive examples, the word "only" is used in different contexts, without a clear specific condition or purpose preceding it.. Example: "All graphs, charts and other visual representations are shown for illustrative purposes only…" Answer:
@@
GPT-4 Failing To Explain Induction-Style Prompts Examples where the attention head is active: out huge amounts of money to go there … of people willing to shell out huge amounts of money to go there How many minutes are there between 4:18 PM … How many minutes are there accustomed - is there … time to become accustomed - is there Explanation: The head is active when the document contains a more complex context or multiple ideas and events...
@@ Appendix E Sample of Explainable Attention Head
Explainable Attention Heads When an attention head has a good explanation, it can be used to predict whether the attention head activates its associated neuron for a given prompt. For example, for an attention head associated with a ’ way’ neuron, if we have the following explanation: Explanation. The attention head in question seems to activate when the phrase "all the way" is used to indicate a complete path, journey, or exhaustive explanation. It is likely looking for contexts where "way" is at the end of a phrase that describes a comprehensive or continuous extent of something. In contrast, the attention head is inactive when "way" is used within idiomatic expressions, comparative phrases (e.g., "the other way around"), or as part of a more metaphorical or abstract concept (e.g., "think of it that way," "put it this way," "stay that way"). The autolabeller successfully classifies the following examples:
Examples where the head is active: FROM 0 TO 5, AND THE Y AXIS MUST INCLUDE THE VALUES FROM 236 ALL THE way stretches as far west as South Dakota, through the northernmost reaches of the U.S. soybean-producing region, all the way bounding from the Oklahoma panhandle in the 1870s all the way Examples where the head is inactive: about why things are the way know G-d’s plans/why He arranges things the way "Wait a minute, it does not work that way why things are the way crosses under the New York State Thruway (Interstate 87). Past the Thruway, the commercial buildings give way
@@
Appendix F More Examples of Explainable Attention Head Activity
F.1 GPT2-Medium
  • (20, 2312, 225), "day"
    • Explanation: Activates with "one day" for future aspirations or developments.
    • Examples: "it may one day", "hoping that will one day"
  • (20, 2312, 66), "day"
    • Explanation: Activates with "present-day" or "modern-day" to describe the current state.
    • Examples: "in modern-day", "was the location in present-day"
  • (21, 2629, 228), "right"
    • Explanation: Activates with "in the right" implying correct direction.
    • Examples: "point me in the right direction", "to point them in the right"
  • (21, 1031, 20), "off"
    • Explanation: Activates with "paid off" or "better off" in contexts of positive outcomes.
    • Examples: "preparation pays off", "hard work and grit always pay off"
  • (21, 3049, 36), "but"
    • Explanation: Activates with "no choice but" to indicate compulsion or lack of alternatives.
    • Examples: "have no choice but to", "no option but to"
  • (23, 2257, 59), "some"
    • Explanation: Activates with "quite some" indicating an unspecified quantity or duration.
    • Examples: "been wanting to check out for quite some time", "planned feature for inclusion for quite some time"
F.2 GPT2-Small
  • (10, 2003, 11), "so"
    • Explanation: Activates with phrases "or so" or "so" to approximate quantity or degree.
    • Examples: "Another 20,000 or so", "Well, a week or so"
  • (10, 1453, 109), "then"
    • Explanation: Activates in shell scripting syntax, particularly with "if" and "then" statements.
    • Examples: "if [ "\$?" -eq "0" ] ; then", "if [[ \${PV} == ’$9$’ ]] ; then"
  • (11, 822, 115), "time"
    • Explanation: Activates with "this time" emphasizing the current instance of a recurring action.
    • Examples: "Austin approaches Grandma again, this time", "is back in the news again — this time"
  • (11, 3015, 10), "on"
    • Explanation: Activates with "and so on" to imply a non-exhaustive list or continuation.
    • Examples: "input/output devices, and so on", "medical images, demographics, and so on"
F.3 Pythia-160m
  • (9, 602, 0), "about"
    • Explanation: Activates in the context of expressing a gap in knowledge or information.
    • Examples: "very little is known about", "however, little is known about"
  • (9, 2251, 71), "all"
    • Explanation: Activates when listing a series followed by "all" as the capper.
    • Examples: "Felipe Anderson, Arthur Masuaku, all", "Hotels.com, Expedia, Hotwire, all"
  • (10, 2509, 80), "more"
    • Explanation: Activates when "more" is preceded by "much" at the end of a list.
    • Examples: "Valkyrie Queen, Maggie and the Martians and so much more", "jeans, bags, skirts, and much more"
  • (10, 2734, 10), "on"
    • Explanation: Activates with lists that suggest continuation, like "and so on".
    • Examples: "be on the bottom of the stack, and so on", "With one citation in superscript, and so on"
F.4 Pythia-1.4b
  • (20, 750, 48), "or"
    • Explanation: Activates in legal disclaimers, licensing terms, or error handling in code.
    • Examples: "mysqli_query(\$dbc,\$query) or", "unless required by applicable law or"
  • (20, 3507, 63), "this"
    • Explanation: Activates when setting up statements of purpose in research.
    • Examples: "The aim of this study", "The purpose of this investigation"
  • (21, 3152, 290), "at"
    • Explanation: Activates in contexts of risk or vulnerability.
    • Examples: "Your capital is at risk", "lives are at stake"
  • (23, 7671, 273), "like"
    • Explanation: Activates to introduce examples in comparisons.
    • Examples: "options such as x and y like", "categories such as a and b like"